# SEEK: Semantic Reasoning for Object Goal Navigation in Real World Inspection Tasks

### v1

@@ V-B Simulation Study
We evaluate SEEK on object-goal navigation tasks in a simulated environment of an office space. We use the ROS Gazebo simulator with the Spot model from Boston Dynamics. In this experiment, the robot is tasked to search and navigate to the nearest fire extinguisher on the first set of simulations and to the coffee mug on the second simulation. The environment consists of $21$ rooms based on the office blueprints (Fig. 4). The placement of fire extinguishers and coffee mugs is based on the real placement in our office. We simulate the object detection and localization using a semantic observation model [2], with model parameters based on empirical data from our hardware experiments in the office. The RSN network is not trained any of the object placement distribution. The entire software stack and simulations were performed on a laptop with an Intel i9 CPU.
We first evaluate the performance of SEEK to find the nearest fire extinguisher across $7$ runs with different starting locations. We compare our approach against the SB2G method that uses geometric coverage to explore the environment until it detects the target object. The SPL performance comparison and a sample of the robot path is presented in Fig. 5 and Fig. 4. Our approach reasons what type of rooms in the office that usually have a fire extinguisher. The semantic-guided global planner successfully directs the robot to the nearest room with a fire extinguisher. Compared to the SB2G method, SEEK significantly reduces the robot distance travel to detect the object. On multiple runs starting from the same location, the distance variation is much lower compared to SB2G, which uses geometric coverage that can stochastically produce different path to maximize space coverage. On various starting locations, our method has a higher SPL than SB2G, validating the benefit of the semantic-guided global planner.
@@ V-C Hardware Results
We validate our approach by deploying SEEK on the Boston Dynamics Spot legged robot for an inspection task within an office building. The robot is equipped with a LIDAR and three cameras for navigation and semantic observation. The robot uses a LIDAR-based SLAM method for localization [43] and a YOLO-based model [42] for object detection. We build the DSG using the office blueprint and use the RSN without any fine-tuning to the office data. The object detection model runs on an NVIDIA Jetson AGX Xavier computer, and the rest of the stack runs on an Intel i7 computer on the robot. The robot receives an object-goal inspection query through a remote user interface. In this experiment, the robot is tasked to locate and go to a fire extinguisher, a commonly inspected object, due to its critical role in safety and emergency response. There are three fire extinguishers located in the office, including in the kitchen area, battery charging room, and the office receptionist.
Using SEEK, the robot efficiently searches and navigates to the nearest fire extinguisher (Fig. 7). The global policy directs the robot to search rooms/regions where fire extinguishers are usually located. In the experiment presented in  Fig. 7, our approach is able to navigate to the closest fire extinguisher with $SPL=0.93$. SEEK reduces the inspection time by directing the robot to search in rooms that have a higher probability of finding fire extinguishers. By distilling the semantic knowledge to the RSN, the robot does not need to access semantic knowledge directly from LLMs or large foundational models that require significant computational resources onboard or access to the internet. The RSN inference and the global planning can be performed in real time, making our approach feasible for real-world object inspection.
@@ B. Planning Evaluation for Target Objects Located in Unexpected Locations
We visualize one example when the object is placed in the least expected location in the PPE storage room (Room number 13). Initially, our model predicts the probability of finding a coffee mug in the room is $0.05$. Fig. 9 shows the sequence of robot’s global action and changes in the probability estimate of finding the object. At the beginning of the run, the robot moves to nearby rooms with higher probability. After searching rooms where a coffee mug is usually located, it visits less common rooms and finds the coffee mug after 25 global planning steps. After finding the target object, the updated probability of finding it in the PPE storage is much higher than in other rooms.

### v2

@@ V-B Simulation Study
We evaluate SEEK on object-goal navigation tasks in a simulated environment of an office space. We use the ROS Gazebo simulator with the Spot model from Boston Dynamics. In this experiment, the robot is tasked to search and navigate to the nearest fire extinguisher on the first set of simulations and to the coffee mug on the second simulation. The environment consists of $21$ rooms based on the office floor plan (Fig. 4). The placement of fire extinguishers and coffee mugs is based on the real placement in our office. We simulate the object detection and localization using a semantic observation model [2], with model parameters based on empirical data from our hardware experiments in the office. The RSN network is not trained in any of the object placement distributions. The entire software stack and simulations were performed on a laptop with an Intel i9 CPU.
We first evaluate the performance of SEEK to find the nearest fire extinguisher across $7$ runs with different starting locations. We compare our approach against the SB2G method that uses geometric coverage to explore the environment until it detects the target object. The SPL performance comparison and a sample of the robot path are presented in Fig. 5 and Fig. 4. Our approach reasons what type of rooms in the office that usually have a fire extinguisher. The semantic-guided global planner successfully directs the robot to the nearest room with a fire extinguisher. Compared to the SB2G method, SEEK significantly reduces the distance traveled by the robot to find the object. On multiple runs starting from the same location, the distance variation is much lower compared to SB2G, which uses geometric coverage that can stochastically produce different paths to maximize space coverage. On various starting locations, our method has a higher SPL than SB2G, validating the benefit of the semantic-guided global planner.
@@ V-C Hardware Results
We validate our approach by deploying SEEK on the Boston Dynamics Spot legged robot for an inspection task within an office building. The robot is equipped with a LIDAR and three cameras for navigation and semantic observation. The robot uses a LIDAR-based SLAM method for localization [43] and a YOLO-based model [42] for object detection. We build the DSG using the office floor plan and use the RSN without any fine-tuning to the office data. The object detection model runs on an NVIDIA Jetson AGX Xavier computer, and the rest of the stack runs on an Intel i7 computer on the robot. The robot receives an object-goal inspection query through a remote user interface. In this experiment, the robot is tasked to locate and go to a fire extinguisher, a commonly inspected object, due to its critical role in safety and emergency response. There are three fire extinguishers located in the office, including in the kitchen area, battery charging room, and the office receptionist.
Using SEEK, the robot efficiently searches and navigates to the nearest fire extinguisher (Fig. 7). The global policy directs the robot to search rooms/regions where fire extinguishers are usually located. In the experiment presented in  Fig. 7, our approach is able to navigate to the closest fire extinguisher with $SPL=0.93$. SEEK reduces the inspection time by directing the robot to search in rooms that have a higher probability of finding fire extinguishers. By distilling the semantic knowledge to the RSN, the robot does not need to access semantic knowledge directly from LLMs or large foundational models that require significant computational resources onboard or access to the internet. The RSN inference and the global planning can be performed in real-time, making our approach feasible for real-world object inspection.
@@ B. Planning Evaluation for Target Objects Located in Unexpected Locations
We visualize one example when the object is placed in the least expected location in the PPE storage room (Room number 13). Initially, our model predicts the probability of finding a coffee mug in the room is $0.05$. Fig. 9 shows the sequence of the robot’s global action and changes in the probability estimate of finding the object. At the beginning of the run, the robot moves to nearby rooms with higher probability. After searching rooms where a coffee mug is usually located, it visits less common rooms and finds the coffee mug after 25 global planning steps. After finding the target object, the updated probability of finding it in the PPE storage is much higher than in other rooms.